Catalina Films produces video shorts using digital editing equipment (K) and editors (L).

 The firm has the production function where Q=K1/5L4/5 where Q is the hours of edited footage. The wage is $160, and the rental rate of capital is $40. The firm wants to produce 3,000 units of output at the lowest possible cost. Find the least costly combination of labor and capital to produce 3000 units 



1
Expert's answer
2021-04-16T07:28:43-0400

The least costly combination of labor and capital to produce 3000 units is possible if MPL/MPK = w/r.

"MPL = Q'(L) = 4\/5(K\/L)^{1\/5},"

"MPK = Q'(K) = 1\/5(L\/K)^{4\/5},"

4K/L = 160/40,

K = L.

"Q=K^{1\/5} K^{4\/5} = 3,000,"

K = L = 3,000.